OpenStudy (anonymous):

find the inverse function f^-1(x) if f(x)= 4(x-2)

OpenStudy (irishboy123):

to make it simpler, call f(x) something like y or z or Bob, so that it just becomes another variable/ placeholder you get y = 4(x-2) then you need to end up with a statement that expresses x as some function of y so divide each side by 4 y/4 = x-2 then add 2 to each side and so on...